# Classical and Quantum Heuristics for the Binary Paint Shop Problem

### v1

@@ I Introduction
The advent of noisy intermediate-scale quantum (NISQ) devices [1, 2, 3] has opened new opportunities toward achieving near-term quantum advantage [4]. Although constrained by limited qubit connectivity, finite coherence times, and the absence of fault-tolerant quantum error correction, such devices can nevertheless execute relatively shallow quantum circuits with reasonable fidelity [5, 6]. These limitations have spurred the development of algorithms designed to maximise the utility of scarce quantum resources, most notably variational quantum algorithms (VQAs) [7, 8, 9, 10], which combine quantum state preparation with classical parameter optimisation to adapt to device imperfections while tackling classically hard problems. A prominent example is the Quantum Approximate Optimisation Algorithm (QAOA) [9], which has been extensively studied for combinatorial optimisation [9, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21] through numerical benchmarks [22, 23, 24], theoretical analyses [25, 26, 27, 28, 29], and numerous algorithmic variants [30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. Beyond optimisation, QAOA is also computationally universal [40, 41] and capable of generating distributions that are classically hard to sample from [42, 43].
@@
One application of particular interest is the Binary Paint Shop Problem (BPSP) [44, 45, 46], an $\mathsf{APX}$-hard optimisation problem arising in automotive manufacturing. The input is a sequence of $2n$ cars comprising $n$ distinct models, each appearing twice. The objective is to paint each car in one of two colours such that the two occurrences of every model receive different colours, while minimising the number of colour transitions between consecutive cars. Performance is quantified by the paint swap ratio—defined as the average number of adjacent pairs of cars painted in different colours—which directly reflects production efficiency and material usage. Streif et al. [47] showed that QAOA at depth $p=7$ (QAOA7) achieves an expected paint swap ratio of $0.393$, surpassing the Recursive Greedy (RG) algorithm [48, 49], which has a proven expected ratio of $0.4$ and was, at the time, the best known classical heuristic for the BPSP. More recently, a new classical heuristic—the Recursive Star Greedy (RSG) heuristic [50]—was proposed, with a conjectured expected ratio of $0.361$. While QAOA7’s ability to outperform RG highlights the potential of quantum algorithms for solving complex real-world problems, matching or exceeding the conjectured performance of RSG would require quantum circuits far deeper than current devices can sustain. Indeed, noise accumulates rapidly with circuit depth [51, 52, 53, 54, 55, 56, 57, 58], and while error mitigation can help [59, 60, 61, 62, 63, 22, 64, 65], its overhead often outweighs the modest performance gains [66, 67].
@@
Motivated by the limitations of deep QAOA circuits, we investigate two state-of-the-art low-depth variants—eXpressive QAOA (XQAOA) [34] and Recursive QAOA (RQAOA) [30]—both of which have outperformed the best known classical algorithms for the maximum cut (MaxCut) problem at $p=1$. We begin by establishing a formal reduction of BPSP to the weighted MaxCut problem and subsequently to an Ising spin model. Central to this reduction is the Initial Car Colour (ICC) encoding, which we formalise based on a technique introduced by Streif et al. [47], that halves the number of decision variables. We further derive exact interaction strengths from BPSP constraints and prove new structural properties of the resulting graphs, including a maximum degree bound and an explicit relation between total graph weight and paint swap cost. Together, these results provide a rigorous foundation for applying QAOA and its variants to BPSP and may be of independent interest for related optimisation problems.
@@
We note that, during the preparation of this manuscript, a related study by Mooney et al. [68] appeared on the arXiv, also applying RQAOA to the BPSP. Their approach differs from ours in that they avoid direct optimisation of QAOA parameters within each recursion. Instead, they leverage parameter concentration [69] and parameter transfer [70], setting QAOA parameters based on values optimised for other instances. Benchmarking RQAOA for $1\leq p\leq 3$ on instances of up to 20 cars, they report no noticeable performance loss under parameter transfer. While these results indicate that transferred parameters can sustain RQAOA performance on small instances, our findings reveal that even with fully optimised parameters, RQAOA suffers diminishing returns as problem size increases. Beyond the shared application of RQAOA to the BPSP, the focus and conclusions of our works are distinct.
@@ II.1 The Binary Paint Shop Problem
The binary paint shop problem (BPSP) [44, 45, 46] is a combinatorial optimisation problem that arises in the paint shops of the automobile industry. A production line receives a fixed sequence of $n$ distinct car models, each appearing exactly twice in the sequence. The order is predetermined by upstream and downstream factory stages—such as the body shop before the paint shop, and trim/chassis/assembly after it—and cannot be rearranged. Each car must be painted in two different colours (e.g., red and blue), but the order in which each car receives its two colours is left to be decided. This assignment, called a colouring, determines how colours are applied across the sequence. The objective is to choose a colouring that minimises the number of colour changes between consecutive cars, since every change requires additional setup time, increases paint waste, and reduces throughput.
@@
The BPSP is $\mathsf{NP}$-complete in its decision form [45] and $\mathsf{APX}$-hard in its optimisation form [44, 46]. In other words, unless $\mathsf{P}=\mathsf{NP}$, no polynomial-time classical algorithm can solve all instances exactly, nor approximate them arbitrarily well via a polynomial-time approximation scheme. Moreover, assuming the Unique Games Conjecture, BPSP is $\mathsf{NP}$-hard to approximate within any constant factor [46]. This is in contrast to problems such as MaxCut, which admit polynomial-time classical algorithms that achieve constant-factor approximations [71].
Given these complexities, several heuristic approaches have been proposed for the BPSP, each producing solutions with colour changes linear in $n$ [72, 49, 48]. The simplest is the red-first algorithm [48], which colours the first occurrence of every car red and the second blue. The greedy algorithm [72, 49] improves on this by selecting at each step the colour that minimises immediate paint changes, switching only when necessary to satisfy the requirement that each car be painted in both colours. The recursive greedy (RG) algorithm [48, 49] offers a further improvement by recursively removing the last-occurring car from the sequence, optimally colouring the reduced instance, and then reinserting the removed pair in the best possible way. More recently, Hančl et al. [50] proposed the recursive star greedy (RSG) algorithm, a refinement of RG that introduces a temporary “undecided” colour (star) for car pairs that can be coloured either way without affecting the current swap count; during recursion, if placing a car next to a starred car would otherwise increase the swap count, both occurrences of that starred car are recoloured with binary colours to avoid the increase, and the final colouring is obtained by replacing any remaining stars with arbitrary legal binary assignments. Detailed explanations of these heuristics are provided in section B.2. The asymptotic performances of red-first, greedy, and RG are proven, with $\mathbb{E}_{\text{RF}}(\Delta_{C})=2n/3$, $\mathbb{E}_{\text{G}}(\Delta_{C})=n/2$, and $\mathbb{E}_{\text{RG}}(\Delta_{C})=2n/5$; in contrast, RSG’s performance is unproven but conjectured to be $\mathbb{E}_{\text{RSG}}(\Delta_{C})=0.361n$, an improvement over RG. Despite these advances, numerical results indicate that the optimal number of colour changes grows linearly in $n$, but with a much smaller constant factor than current heuristics achieve. A recent lower bound of $\mathbb{E}_{\text{OPT}}(\Delta_{C})\geq 0.214n-o(n)$ by Hančl et al. [50] confirms that existing heuristics remain far from optimal in the asymptotic limit, leaving room for quantum algorithms to yield improvements.
@@
We conclude this subsection with an illustrative example of the BPSP, shown in fig. 1 with the corresponding results summarised in table 1. The example compares the four classical heuristics described earlier on the same instance. As expected from their asymptotic behaviour, red-first produces the highest number of swaps, while RSG achieves the fewest and, in this case, matches the optimal solution. It is worth noting that although RG generally outperforms the standard greedy algorithm—as it does in this example—there are instances where the greedy algorithm produces fewer paint swaps [49].
@@ II.2 Quantum Approximate Optimisation Algorithm
The Quantum Approximate Optimisation Algorithm (QAOA) [9] is a variational quantum algorithm for approximating solutions to combinatorial optimisation problems. The problem is encoded into a problem Hamiltonian $H_{P}$, whose ground state represents the optimal solution. QAOA seeks to approximate this ground state by preparing a depth-$p$ variational quantum circuit, parameterised by $2p$ real angles, and optimising them to minimise the expectation value $\left\langle H_{P}\right\rangle$. Measuring the resulting quantum state in the computational basis produces candidate bit strings, from which a low-cost solution is selected.
@@ II.3 Expressive QAOA
The Expressive QAOA (XQAOA) [34] is an overparameterised extension of standard QAOA, designed to increase ansatz expressiveness at shallow circuit depths. While standard QAOA uses a single $\gamma$ for all terms in $H_{P}$ and a single $\beta$ for all terms in $H_{M},\mathrm{XQAOA}$ assigns independent parameters to each term of both Hamiltonians. This allows different qubits and couplings to be rotated by distinct angles, enabling a richer exploration of the solution space.
@@
The XQAOA ansatz admits four notable configurations, distinguished by the constraints placed on the mixing unitary angles. The Multi-Angle QAOA (MA-QAOA) [33] is obtained by setting all $\alpha_{j}=0$, yielding a multi-angle extension of the standard X-only QAOA mixer. The XY mixer is the most general form, allowing independent $\alpha_{j}$ and $\beta_{j}$ for each qubit. The Y mixer restricts the ansatz to $Y$-rotations by setting all $\beta_{j}=0$, while the X=Y mixer applies both $X$ and $Y$ rotations with a shared angle on each qubit, enforcing $\alpha_{j}=\beta_{j}$. Numerical results indicate that the X=Y mixer outperforms the other three configurations; therefore, all references to XQAOA in this work pertain to the X=Y mixer.
@@ II.4 Recursive QAOA
The Recursive QAOA (RQAOA) [30] is a non-local variant of the QAOA designed to enhance QAOA’s performance, especially at shallow circuit depths. RQAOA leverages QAOA as a subroutine to recursively reduce the problem size until it becomes tractable for brute-force methods. Specifically, RQAOA1—RQAOA that uses QAOA1 as its subroutine—has been shown to outperform state-of-the-art classical algorithms for MaxCut and similar problems [30, 73, 74]. Additionally, RQAOA1 provides performance guarantees for particular graph structures [30, 75, 76] and has been successfully applied to small-scale, real-world problems [77, 78].
@@ III Recasting BPSP as an Ising Model
We now present a complete and provably equivalent mapping of the BPSP to an Ising model suitable for quantum optimisation. Our approach formalises the Initial-Car-Colour (ICC) encoding, inspired by Streif et al. [47], which reduces the number of decision variables by half. Using this encoding we construct, for any BPSP instance, a weighted graph whose MaxCut value directly corresponds to the optimal paint swap cost. In particular, we establish a polynomial-time reduction from BPSP to weighted MaxCut, which in turn, provides an exact Ising Hamiltonian whose ground state encodes the optimal solution. The full technical development, with proofs and derivations, is given in the appendix; here we present only the essential ingredients needed for the remainder of the paper.
@@ III.2 BPSP Graph Construction
The BPSP graph has two particularly useful structural properties. First, we show in proposition 8 that its maximum degree is bounded by four, since each car can be adjacent to at most two distinct neighbours across its two occurrences. Second, as shown by Streif et al. [47], the edge weights take values in $\{\pm 1,-2\}$. As $n\rightarrow\infty$, edges of weight -2 vanish in probability, while the remaining couplings occur with limiting probabilities: -1 with probability $2/3$ and +1 with probability $1/3$.
@@ IV Numerical Results
In this section, we benchmark the performance of XQAOA1 and RQAOA1 on the Binary Paint Shop Problem, comparing them against the known classical heuristics: Red-First, Greedy, Recursive Greedy, and Recursive-Star Greedy. All BPSP instances were generated randomly using a quantum random number generator [79, 80]. Since both XQAOA1 and QAOA1 admit closed-form expressions for their expectation values—given respectively by Theorem 3, Eq. (22), and Theorem 1, Eq. (13) of [34]—all benchmarks were performed entirely on a classical computer, without the use of quantum hardware. For XQAOA1, the large number of parameters motivated the use of the JAX framework for efficient gradient computation via backpropagation, with optimisation carried out using the Limited-Memory Broyden–Fletcher–Goldfarb–Shanno (LBFGS) algorithm [81]. To address the highly non-convex optimisation landscape, each instance was run with 100 random initialisations. For RQAOA1, the pairwise expectation values required at each recursion step were computed analytically as functions of $(\beta,\gamma)$, and the variational parameters were set using the provably near-optimal values obtained by the method of Vijendran et al. [82] [^1] In brief, this method applies to Ising models with arbitrary rational weights, reducing the two-dimensional optimisation over $(\beta,\gamma)$ to a one-dimensional search in $\gamma$, using the Nyquist–Shannon theorem to set the sampling resolution, and a subdivision algorithm to locate the optimal angle in linear time. , eliminating the need for multiple restarts; hence, each instance was simulated only once. The deterministic classical heuristics were likewise run once per instance.

### v2

@@ I Introduction
The advent of noisy intermediate-scale quantum (NISQ) devices [preskill2018quantum, preskill2021quantum, cheng2023noisy] has opened new opportunities toward achieving near-term quantum advantage [huang2025vast]. Although constrained by limited qubit connectivity, finite coherence times, and the absence of fault-tolerant quantum error correction, such devices can nevertheless execute relatively shallow quantum circuits with reasonable fidelity [beverland2022assessing, lau2022nisq]. These limitations have spurred the development of algorithms designed to maximise the utility of scarce quantum resources, most notably variational quantum algorithms (VQAs) [cerezo2021variational, bharti2022noisy, farhi2014quantum, peruzzo2014variational], which combine quantum state preparation with classical parameter optimisation to adapt to device imperfections while tackling classically hard problems. A prominent example is the Quantum Approximate Optimisation Algorithm (QAOA) [farhi2014quantum], which has been extensively studied for combinatorial optimisation [farhi2014quantum, farhi2015quantum, vikstaal2020applying, harrigan2021quantum, farhi2022quantum, anschuetz2019variational, harwood2021formulating, amaro2022case, azad2022solving, fitzek2022applying, fitzek2024applying, ruan2023quantum] through numerical benchmarks [zhou2020quantum, lotshaw2021empirical, golden2023numerical], theoretical analyses [bittel2021training, hadfield2022analytical, wurtz2021maxcut, ng2024analytical, benchasattabuse2025lower], and numerous algorithmic variants [bravyi2020obstacles, yu2022quantum, chandarana2022digitized, herrman2022multi, vijendran2023expressive, gaidai2024performance, yoshioka2023fermionic, bartschi2020grover, zhu2022multi, blekos2024review]. Beyond optimisation, QAOA is also computationally universal [lloyd2018quantum, morales2020universality] and capable of generating distributions that are classically hard to sample from [farhi2016quantum, dalzell2020how].
@@
One application of particular interest is the Binary Paint Shop Problem (BPSP) [epping2001some, epping2004complexity, gupta2013approximability], an $\mathsf{APX}$-hard optimisation problem arising in automotive manufacturing. The input is a sequence of $2n$ cars comprising $n$ distinct models, each appearing twice. The objective is to paint each car in one of two colours such that the two occurrences of every model receive different colours, while minimising the number of colour transitions between consecutive cars. Performance is quantified by the paint swap ratio—defined as the average number of adjacent pairs of cars painted in different colours—which directly reflects production efficiency and material usage. Streif et al. [streif2021beating] showed that QAOA at depth $p=7$ (QAOA7) achieves an expected paint swap ratio of $0.393$, surpassing the Recursive Greedy (RG) algorithm [andres2011some, andres2021greedy], which has a proven expected ratio of $0.4$ and was, at the time, the best known classical heuristic for the BPSP. More recently, a new classical heuristic—the Recursive Star Greedy (RSG) heuristic [hanvcl2023improved]—was proposed, with a conjectured expected ratio of $0.361$. While QAOA7’s ability to outperform RG highlights the potential of quantum algorithms for solving complex real-world problems, matching or exceeding the conjectured performance of RSG would require quantum circuits far deeper than current devices can sustain. Indeed, noise accumulates rapidly with circuit depth [xue2021effects, wang2021noise, marshall2020characterizing, alam2019analysis, alam2020design, streif2021quantum, anschuetz2022quantum, stilck2021limitations], and while error mitigation can help [weidinger2023error, shaydulin2021error, botelho2022error, koh2024readout, kandala2019error, zhou2020quantum, koczor2021exponential, leymann2020bitter], its overhead often outweighs the modest performance gains [takagi2022fundamental, quek2024exponentially].
@@
Motivated by the limitations of deep QAOA circuits, we investigate two state-of-the-art low-depth variants—eXpressive QAOA (XQAOA) [vijendran2023expressive] and Recursive QAOA (RQAOA) [bravyi2020obstacles]—both of which have outperformed the best known classical algorithms for the maximum cut (MaxCut) problem at $p=1$. We begin by establishing a formal reduction of BPSP to the weighted MaxCut problem and subsequently to an Ising spin model. Central to this reduction is the Initial Car Colour (ICC) encoding, which we formalise based on a technique introduced by Streif et al. [streif2021beating], that halves the number of decision variables. We further derive exact interaction strengths from BPSP constraints and prove new structural properties of the resulting graphs, including a maximum degree bound and an explicit relation between total graph weight and paint swap cost. Together, these results provide a rigorous foundation for applying QAOA and its variants to BPSP and may be of independent interest for related optimisation problems.
@@
We note that, during the preparation of this manuscript, a related study by Mooney et al. [mooney2025optimization] appeared on the arXiv, also applying RQAOA to the BPSP. Their approach differs from ours in that they avoid direct optimisation of QAOA parameters within each recursion. Instead, they leverage parameter concentration [akshay2021parameter] and parameter transfer [shaydulin2023parameter], setting QAOA parameters based on values optimised for other instances. Benchmarking RQAOA for $1\leq p\leq 3$ on instances of up to 20 cars, they report no noticeable performance loss under parameter transfer. While these results indicate that transferred parameters can sustain RQAOA performance on small instances, our findings reveal that even with fully optimised parameters, RQAOA suffers diminishing returns as problem size increases. Beyond the shared application of RQAOA to the BPSP, the focus and conclusions of our works are distinct.
@@ II.1 The Binary Paint Shop Problem
The binary paint shop problem (BPSP) [epping2001some, epping2004complexity, gupta2013approximability] is a combinatorial optimisation problem that arises in the paint shops of the automobile industry. A production line receives a fixed sequence of $n$ distinct car models, each appearing exactly twice in the sequence. The order is predetermined by upstream and downstream factory stages—such as the body shop before the paint shop, and trim/chassis/assembly after it—and cannot be rearranged. Each car must be painted in two different colours (e.g., red and blue), but the order in which each car receives its two colours is left to be decided. This assignment, called a colouring, determines how colours are applied across the sequence. The objective is to choose a colouring that minimises the number of colour changes between consecutive cars, since every change requires additional setup time, increases paint waste, and reduces throughput.
@@
The BPSP is $\mathsf{NP}$-complete in its decision form [epping2004complexity] and $\mathsf{APX}$-hard in its optimisation form [epping2001some, gupta2013approximability]. In other words, unless $\mathsf{P}=\mathsf{NP}$, no polynomial-time classical algorithm can solve all instances exactly, nor approximate them arbitrarily well via a polynomial-time approximation scheme. Moreover, assuming the Unique Games Conjecture, BPSP is $\mathsf{NP}$-hard to approximate within any constant factor [gupta2013approximability]. This is in contrast to problems such as MaxCut, which admit polynomial-time classical algorithms that achieve constant-factor approximations [goemans1995improved].
Given these complexities, several heuristic approaches have been proposed for the BPSP, each producing solutions with colour changes linear in $n$ [amini2010greedy, andres2021greedy, andres2011some]. The simplest is the red-first algorithm [andres2011some], which colours the first occurrence of every car red and the second blue. The greedy algorithm [amini2010greedy, andres2021greedy] improves on this by selecting at each step the colour that minimises immediate paint changes, switching only when necessary to satisfy the requirement that each car be painted in both colours. The recursive greedy (RG) algorithm [andres2011some, andres2021greedy] offers a further improvement by recursively removing the last-occurring car from the sequence, optimally colouring the reduced instance, and then reinserting the removed pair in the best possible way. More recently, Hančl et al. [hanvcl2023improved] proposed the recursive star greedy (RSG) algorithm, a refinement of RG that introduces a temporary “undecided” colour (star) for car pairs that can be coloured either way without affecting the current swap count; during recursion, if placing a car next to a starred car would otherwise increase the swap count, both occurrences of that starred car are recoloured with binary colours to avoid the increase, and the final colouring is obtained by replacing any remaining stars with arbitrary legal binary assignments. Detailed explanations of these heuristics are provided in section B.2. The asymptotic performances of red-first, greedy, and RG are proven, with $\mathbb{E}_{\text{RF}}(\Delta_{C})=2n/3$, $\mathbb{E}_{\text{G}}(\Delta_{C})=n/2$, and $\mathbb{E}_{\text{RG}}(\Delta_{C})=2n/5$; in contrast, RSG’s performance is unproven but conjectured to be $\mathbb{E}_{\text{RSG}}(\Delta_{C})=0.361n$, an improvement over RG. Despite these advances, numerical results indicate that the optimal number of colour changes grows linearly in $n$, but with a much smaller constant factor than current heuristics achieve. A recent lower bound of $\mathbb{E}_{\text{OPT}}(\Delta_{C})\geq 0.214n-o(n)$ by Hančl et al. [hanvcl2023improved] confirms that existing heuristics remain far from optimal in the asymptotic limit, leaving room for quantum algorithms to yield improvements.
@@
We conclude this subsection with an illustrative example of the BPSP, shown in fig. 1 with the corresponding results summarised in table 1. The example compares the four classical heuristics described earlier on the same instance. As expected from their asymptotic behaviour, red-first produces the highest number of swaps, while RSG achieves the fewest and, in this case, matches the optimal solution. It is worth noting that although RG generally outperforms the standard greedy algorithm—as it does in this example—there are instances where the greedy algorithm produces fewer paint swaps [andres2021greedy].
@@ II.2 Quantum Approximate Optimisation Algorithm
The Quantum Approximate Optimisation Algorithm (QAOA) [farhi2014quantum] is a variational quantum algorithm for approximating solutions to combinatorial optimisation problems. The problem is encoded into a problem Hamiltonian $H_{P}$, whose ground state represents the optimal solution. QAOA seeks to approximate this ground state by preparing a depth-$p$ variational quantum circuit, parameterised by $2p$ real angles, and optimising them to minimise the expectation value $\left\langle H_{P}\right\rangle$. Measuring the resulting quantum state in the computational basis produces candidate bit strings, from which a low-cost solution is selected.
@@ II.3 Expressive QAOA
The Expressive QAOA (XQAOA) [vijendran2023expressive] is an overparameterised extension of standard QAOA, designed to increase ansatz expressiveness at shallow circuit depths. While standard QAOA uses a single $\gamma$ for all terms in $H_{P}$ and a single $\beta$ for all terms in $H_{M},\mathrm{XQAOA}$ assigns independent parameters to each term of both Hamiltonians. This allows different qubits and couplings to be rotated by distinct angles, enabling a richer exploration of the solution space.
@@
The XQAOA ansatz admits four notable configurations, distinguished by the constraints placed on the mixing unitary angles. The Multi-Angle QAOA (MA-QAOA) [herrman2022multi] is obtained by setting all $\alpha_{j}=0$, yielding a multi-angle extension of the standard X-only QAOA mixer. The XY mixer is the most general form, allowing independent $\alpha_{j}$ and $\beta_{j}$ for each qubit. The Y mixer restricts the ansatz to $Y$-rotations by setting all $\beta_{j}=0$, while the X=Y mixer applies both $X$ and $Y$ rotations with a shared angle on each qubit, enforcing $\alpha_{j}=\beta_{j}$. Numerical results indicate that the X=Y mixer outperforms the other three configurations; therefore, all references to XQAOA in this work pertain to the X=Y mixer.
@@ II.4 Recursive QAOA
The Recursive QAOA (RQAOA) [bravyi2020obstacles] is a non-local variant of the QAOA designed to enhance QAOA’s performance, especially at shallow circuit depths. RQAOA leverages QAOA as a subroutine to recursively reduce the problem size until it becomes tractable for brute-force methods. Specifically, RQAOA1—RQAOA that uses QAOA1 as its subroutine—has been shown to outperform state-of-the-art classical algorithms for MaxCut and similar problems [bravyi2020obstacles, bravyi2021classical, bravyi2022hybrid]. Additionally, RQAOA1 provides performance guarantees for particular graph structures [bravyi2020obstacles, bae2024recursive, bae2024modified] and has been successfully applied to small-scale, real-world problems [gulbahar2024majority, palackal2023quantum].
@@ III Recasting BPSP as an Ising Model
We now present a complete and provably equivalent mapping of the BPSP to an Ising model suitable for quantum optimisation. Our approach formalises the Initial-Car-Colour (ICC) encoding, inspired by Streif et al. [streif2021beating], which reduces the number of decision variables by half. Using this encoding we construct, for any BPSP instance, a weighted graph whose MaxCut value directly corresponds to the optimal paint swap cost. In particular, we establish a polynomial-time reduction from BPSP to weighted MaxCut, which in turn, provides an exact Ising Hamiltonian whose ground state encodes the optimal solution. The full technical development, with proofs and derivations, is given in the appendix; here we present only the essential ingredients needed for the remainder of the paper.
@@ III.2 BPSP Graph Construction
The BPSP graph has two particularly useful structural properties. First, we show in proposition 8 that its maximum degree is bounded by four, since each car can be adjacent to at most two distinct neighbours across its two occurrences. Second, as shown by Streif et al. [streif2021beating], the edge weights take values in $\{\pm 1,-2\}$. As $n\rightarrow\infty$, edges of weight -2 vanish in probability, while the remaining couplings occur with limiting probabilities: -1 with probability $2/3$ and +1 with probability $1/3$.
@@ IV Numerical Results
In this section, we benchmark the performance of XQAOA1 and RQAOA1 on the Binary Paint Shop Problem, comparing them against the known classical heuristics: Red-First, Greedy, Recursive Greedy, and Recursive-Star Greedy. All BPSP instances were generated randomly using a quantum random number generator [symul2011real, haw2015maximization]. Since both XQAOA1 and QAOA1 admit closed-form expressions for their expectation values—given respectively by Theorem 3, Eq. (22), and Theorem 1, Eq. (13) of [vijendran2023expressive]—all benchmarks were performed entirely on a classical computer, without the use of quantum hardware. For XQAOA1, the large number of parameters motivated the use of the JAX framework for efficient gradient computation via backpropagation, with optimisation carried out using the Limited-Memory Broyden–Fletcher–Goldfarb–Shanno (LBFGS) algorithm [liu1989limited]. To address the highly non-convex optimisation landscape, each instance was run with 100 random initialisations. For RQAOA1, the pairwise expectation values required at each recursion step were computed analytically as functions of $(\beta,\gamma)$, and the variational parameters were set using the provably near-optimal values obtained by the method of Vijendran et al. [vijendran2025near] [^1] In brief, this method applies to Ising models with arbitrary rational weights, reducing the two-dimensional optimisation over $(\beta,\gamma)$ to a one-dimensional search in $\gamma$, using the Nyquist–Shannon theorem to set the sampling resolution, and a subdivision algorithm to locate the optimal angle in linear time. , eliminating the need for multiple restarts; hence, each instance was simulated only once. The deterministic classical heuristics were likewise run once per instance.